OpenStudy (mackenzie2013):

A board that is 2/3 of a yard long needs to be cut into four equal pieces. What will the length of each piece be?

OpenStudy (anonymous):

We need to divide the length by 4, so: \[(2/3)/4\] Dividing by 4 is the same as multiplying by 1/4, so: \[\frac{ 2 }{ 3 } \times \frac{ 1 }{ 4 } = \frac{ 2 }{ 12 } = \frac{ 1 }{ 6 }\]